OpenStudy (anonymous):

f(x) = square root x+1 g(x) = x^2 - x what does (f-g)(8) = ??

OpenStudy (anonymous):

I believe what you mean is \(f(x)=\sqrt{x+1}\) and \(g(x)=x^2-x\). Find \((f\circ g)(8)\). If this is the case the general answer is \((f\circ g)(x)=\sqrt{x^2-x+1}\). From this we see that \((f\circ g)(8)=\sqrt{8^2-8+1}=\sqrt{64-8+1}=\sqrt{57}\).

hero (hero):

^wrong notation

hero (hero):

(f-g)(x) = f(x) - g(x)

OpenStudy (anonymous):

If that's the case the answer would be \(\sqrt{9}-(64-8)=3-56=-53\)